# Low-hallucination Synthetic Captions for Large-Scale Vision-Language Model Pre-training

### v1

@@ Abstract
In recent years, the field of vision-language model pre-training has experienced rapid advancements, driven primarily by the continuous enhancement of textual capabilities in large language models. However, existing training paradigms for multimodal large language models heavily rely on high-quality image-text pairs. As models and data scales grow exponentially, the availability of such meticulously curated data has become increasingly scarce and saturated, thereby severely limiting further advancements in this domain. This study investigates scalable caption generation techniques for vision-language model pre-training and demonstrates that large-scale low-hallucination synthetic captions can serve dual purposes: 1) acting as a viable alternative to real-world data for pre-training paradigms and 2) achieving superior performance enhancement when integrated into vision-language models through empirical validation. This paper presents three key contributions: 1) a novel pipeline for generating high-quality, low-hallucination, and knowledge-rich synthetic captions. Our continuous DPO methodology yields remarkable results in reducing hallucinations. Specifically, the non-hallucination caption rate on a held-out test set increases from $48.2\%$ to $77.9\%$ for a 7B-size model. 2) Comprehensive empirical validation reveals that our synthetic captions confer superior pre-training advantages over their counterparts. Across 15 vision language tasks, the model trained with our data achieves a significant performance gain of at least $6.2\%$ compared to alt-text pairs and other previous work. In 20 common cognitive domains, the model trained with our data outperforms other counterparts by at least $6.8\%$. Meanwhile, it also offers considerable support in the text-to-image domain. With the support of our dataset, the FID score is reduced by $17.1$ on a real-world validation benchmark and $13.3$ on the MSCOCO validation benchmark. 3) We will release Hunyuan-Recap100M, a low-hallucination and knowledge-intensive synthetic caption dataset, which will enable breakthroughs in vision language pre-training, cross-modal generation, and related applications.
@@ 1 Introduction
With the enormous power of large language models, remarkable performance gains have recently been achieved in a variety of tasks in natural language processing (NLP), computer vision (CV), and also in cross-modal fields (Brown et al., 2020; Chung et al., 2022; Chowdhery et al., 2022; Touvron et al., 2023a; Dosovitskiy et al., 2021; Alayrac et al., 2022; Liu et al., 2023; Achiam et al., 2023; Wang et al., 2024b; Hurst et al., 2024). The Large Language Model (LLM) and the Vision Language Model (VLM) are usually equipped with Transformer (Vaswani et al., 2017) as the backbone and then pre-trained with a tremendous amount of unlabeled data. The strong representation ability of the model, the massive amount of data, and the effective means of training make the foundation models powerful for successfully solving the tasks of vision and language.
@@
While text data has experienced exponential growth and fuels large language models, vision language model pre-training confronts a unique data bottleneck: high-quality text-image alignments remain scarce despite their critical role in training vision language models. Most traditional approaches to building large-scale multimodal datasets mainly rely on Web-based alt-text extraction (Sharma et al., 2018; Changpinyo et al., 2021; Ordonez et al., 2011; Schuhmann et al., 2021, 2022; Byeon et al., 2022; Gadre et al., 2023; Liu et al., 2022; Gu et al., 2022). The web-based alt-text data typically undergoes a lengthy processing pipeline, involving stages such as text quality cleaning, image low-quality filtering, and cross-modal relevance validation, ultimately producing image-text pairs. Although significant data loss occurs at each processing stage and even with these efforts, the final output still suffers from inherent limitations including a low text informativeness density and partial relevance deficiencies between visual content and the corresponding descriptions. As the naturally existing image-text pair data is exhausted, and the quality of data produced by traditional image-text extraction methods drops significantly as the scale expands, the marginal effect they generate in the training of VLMs greatly diminishes. Therefore, new methods are needed to synthesize large-scale, high-quality multimodal data.
@@
Recently, several exploratory investigations have been carried out in multi-modal synthetic data, which have, to some extent, demonstrated their positive impact on various cross-modal tasks (Li et al., 2023a; Betker et al., 2023; Yu et al., 2024; Li et al., 2024b). However, there are significant defects in all of them, mainly reflected in the presence of many hallucinations or a relatively low level of information content. As shown in Figure 1, there are a large number of hallucinations (labeled in red) in the previous synthetic captions, and the captions are rather monotonous, lacking informative expressions. Therefore, in this paper, we focus on addressing two issues: 1. How to establish an effective pipeline for generating low-hallucination synthetic captions? 2. How can more valuable information be injected into synthetic captions to increase the information density of synthetic data?
@@
How to establish an effective pipeline for generating low-hallucination synthetic captions? Large-scale synthetic captions inherently depend on robust tools, and vision language models emerge as a natural and fitting choice. However, hallucination is an inherent problem in the multimodal large language model (Bai et al., 2024). As shown in Figure 1, the data in the Recap-DataComp-1B (Li et al., 2024b) dataset contains a significant amount of hallucinatory information. We further demonstrate the harmfulness of a large amount of hallucinatory data in Section 4. The inevitable and most crucial issue in caption synthesis lies in resolving the hallucination problem. Existing work (Bai et al., 2024) has discussed the causes of hallucinations from the perspectives of data, model, training, and inference. During generation, as the length of the sequence grows, self-attention will focus more on previously generated text tokens, i.e., the attention to visual content is diluted(Wang et al., 2023b, a, 2024a; Huang et al., 2024). We agree that hallucinations in VLMs are mainly caused by uncertainty in decoding visual-related intentions. Therefore, we propose the use of Direct Preference Optimization (DPO) method to eliminate hallucinations in synthesized captions by increasing the confidences of real visual contents and suppressing the confidences of hallucinatory visual contents. Fortunately, this method has proven to be highly effective. Furthermore, in the synthetic caption task, we reveal that DPO adheres to a significant scaling law within a specific data size range. Upon saturation, resampling can further elevate the quality of synthetic captions. As presented in Table 1, when applied to the same dataset, our method significantly enhances the non-hallucination rate, elevating it from 29.7% to 73.3%, in contrast to the approach used in Recap-DataComp-1B (Li et al., 2024b). We elaborate on our hallucination evaluation metrics in Section 4.1 and the continuous DPO method in Section 3.
@@
Table 1: A comparison between existing synthetic caption datasets and our proposed dataset in terms of the number of hallucinations and the hallucination rate. Here, "Length" represents the average length of the caption, "Details" represents the average number of visual details that appear in each caption, "None Hallucinations" indicates the proportion of non-hallucinatory data in the overall data, and "hallucination rate" represents the proportion of hallucinations among all details. A total of 1,000 data samples were taken, and our method reports the corresponding results on the respective data with our synthetic caption pipeline respectively.
How can more valuable information be injected into synthetic captions to increase the information density of synthetic data? The answer to this question can be rather straightforward: infuse as much knowledge as possible into the caption via a well-trained VLM. The injection of knowledge into synthetic captions can be derived from two primary sources. The first source is external text, such as alt-text, that is relevant to the image, while the second source stems from world knowledge encapsulated within a VLM. Usually, a VLM model tends to provide captions that are vague but correct. Therefore, we need to train a VLM model so that it can fully integrate the knowledge in text and its own world knowledge, ultimately generating synthetic captions with high informational value. Specifically, in our methodology, we employ a powerful VLM, namely GPT-4o, in conjunction with a meticulously crafted prompt that encompasses detailed requirements. This combination is utilized to generate labeled data, with a particular emphasis on extracting and focusing on the key information embedded within alt-text. For instance, we focus more on the name of a specific entity, using oxalic acid as an illustration here. After that, we perform manually post-processing on obvious erroneous data to ensure that the labeled data is accurate, contains the key information, and excludes the noise in the injected information. Finally, we train an open source VLM model, specifically Qwen2-VL-7B, using the training data. For details of the prompt and data processing, see Section 3.
By comprehensively leveraging the aforementioned methods, we train a recaption model and utilize it to generate 100 million pieces of data. Our model has achieved a notable enhancement in the evaluation metrics for low-hallucination quantification. Moreover, we carried out a verification of the data’s validity. The results show that vision-language models pre-trained with Hunyuan-Recap100M exhibit substantially improved overall performance compared to those using alt-text data or conventional synthetic captions, particularly showing enhanced capabilities in perception tasks.
Our contributions are as follows.
(1) We tackle the challenge of constructing a synthetic caption pipeline capable of generating captions that are both low in hallucination and rich in knowledge, thereby alleviating the scarcity of high-quality multimodal data essential for vision-language model pre-training.
(2) We propose a series of effective methods, knowledge-enriching SFT and continuous DPO, to eliminate hallucinations and inject knowledge into synthetic captions. Eventually, we trained a recaption model, whose proportion of hallucination-free results far exceeds that of the existing best models. The proportion of hallucination-free results has increased from $48.2\%$ to $77.9\%$. It also exhibits favorable performance in knowledge diversity, as reflected in a more accurate cognition of real-world entities.
(3) We will release a high-quality synthetic caption dataset comprising $100M$ image-text pairs, Hunyuan-Recap100M, with the aim of fostering further advancements within the research community in areas like vision language model pre-training, text-to-image, and text-to-video generation.
@@ 2 Related Work
Vision-Language Model. Vision Language Models (VLMs) can be classified into two distinct categories. The first category consists of models that rely on dual encoders, specifically a vision encoder and a text encoder. These models are typically trained in a contrastive fashion and exhibit remarkable proficiency in tasks such as image recognition. Prominent examples within this group include CLIP (Radford et al., 2021), FLIP (Li et al., 2023b) and SigLIP (Zhai et al., 2023b). The second category of methods focuses on aligning the visual modality with a Large Language Model (LLM). With the rapid development of LLM, the astonishing success of GPT (Brown et al., 2020; OpenAI, 2023), LLaMA (Touvron et al., 2023a, b; Grattafiori et al., 2024), and DeepSeek-LLM (Liu et al., 2024; Guo et al., 2025) has made the empowerment of large language models with visual perception an important and popular research direction. Flamingo (Alayrac et al., 2022) is the first large-scale vision language model, which aligns the visual encoder with a large language model. LLaVA (Liu et al., 2023) presents an instruction tuning methodology for training large multimodal models, thus charting the course for VLM training. Subsequent efforts have been made to refine the details of each module and continuously optimize the overall performance of the model (Wang et al., 2024b; Wu et al., 2024; McKinzie et al., 2024; Chen et al., 2024b; Liu et al., 2024). Although, compared with former multimodal models (Zeng et al., 2021; Singh et al., 2021; Diao et al., 2022; Wang et al., 2022; Zhang et al., 2023a), current LLM based vision language models are much more powerful in terms of capabilities, there has not been much improvement in data utilization; they still rely on large-scale image-text pairs. Similarly, VLMs with dual encoders, as referred to in the first category above, also require large-scale image-text pair data. However, the data of the image-text pairs are gradually being exhausted (Villalobos et al., 2022). The total amount of a large number of cleaned image-text pairs is fixed, and the expansion of data scale can hardly keep up with the requirements of model scaling. Therefore, despite the remarkable advancements achieved in the VLM training paradigm, the utilization of large-scale image-text pairs data has encountered certain limitations. To overcome this issue, a potential solution lies in synthetic data.
Synthetic data. Synthetic data has been extensively utilized in large language models, yielding remarkable results (Zhang et al., 2024; Shao et al., 2024; Zhu et al., 2024; Yang et al., 2024). In addition, there is a growing presence of synthetic data within the multimodal domain. BLIP-2 (Li et al., 2023a) employs an image caption model to synthesize short captions to serve as substitutes for alt-text, and CapFusion (Yu et al., 2024) further enhances the quality of such data using large language models. However, the captions generated through this type of work tend to be rather simplistic, merely representing minor refinements of the original alt-text without adding substantial information value. Recap-DataComp-1B (Li et al., 2024b) aims to transform alt-text using vision language models, yet it encounters the issue that existing VLMs can produce severe hallucinations. Other works focused on dense captioning, such as Allava (Chen et al., 2024a) and Dense-Fusion-1M (Li et al., 2024c), are incapable of injecting rich knowledge in a targeted fashion and, on this basis, maintain relatively low levels of hallucination. In summary, knowledge injection and elimination of hallucinations are two key challenges in the realm of multimodal synthetic data. A significant amount of research has been dedicated to exploring the hallucination problem in VLMs, including aspects such as the causes underlying hallucinations and strategies for their elimination (Bai et al., 2024). However, a comprehensive method that can simultaneously ensure content diversity and maximally eliminate hallucinations in the synthetic visual caption task remains elusive.
Our approach has achieved substantial advances in both hallucination elimination and knowledge injection. Ultimately, its efficacy has been demonstrated in the training of VLMs, establishing our caption dataset as a viable alternative for the large-scale training of vision language models. This eliminates the constraints imposed by the data scale and avoids the challenges associated with data quality degradation resulting from the expansion of the data scale.
@@ 3.1 Continuous DPO
Reinforcement learning has been reported to be an effective method of eliminating hallucinations in Large Language Models (LLMs) (Zhang et al., 2023b; Huang et al., 2025). Direct Preference Optimization (DPO) (Rafailov et al., 2023) emerges as a reinforcement learning based fine-tuning paradigm for LLM that directly optimizes preference data, effectively mitigating harmful content generation, such as hallucinations. Consequently, the application of DPO to mitigate hallucinations in image captions appears to be a natural and logical approach. The optimization objectives of the DPO algorithm are shown in Equation 1 and Equation 2. As noted in Rafailov et al. (2023), the loss function gradient $\mathcal{L}_{\text{DPO}}$ increases the probability of preferred completions $y_{w}$ while simultaneously reducing the probability of dispreferred completions $y_{l}$. Intuitively, by maximizing the probability of samples free of hallucinations and minimizing the probability of samples containing hallucinations, our recaption model can generate captions that are as free from uncertainty and potential hallucinations as possible.
@@
However, in our task, the DPO algorithm does not demonstrate continuous performance improvement with respect to scale. Specifically, after initial rapid scaling up of performance, its performance plateaus, as illustrated in Figure 4. We turn our attention back to Equation2. Here, the examples are weighted according to how much higher the implicit reward model $\hat{r}_{\theta}(x,y)$ rates the dispreferred completions. As we continuously expand the scale of the DPO data, the implicit reward $\hat{r}_{\theta}(x,y)$, which serves as the optimization weight, does not change significantly. Consequently, after reaching a local optimum, the model ceases to optimize further. Therefore, we make slight modifications to the DPO algorithm and propose Continuous DPO (CDPO). Once the performance reaches a bottleneck phase, we switch to a different reference model, resample data using the current model, adjust the implicit reward model $\hat{r}_{\theta}(x,y)$ accordingly, and then re-optimize $\pi_{\theta}$. Furthermore, when CDPO training with preference data is performed, in order to prevent the sample length from serving as a learning signal (e.g., where longer captions may contain more hallucinations), we conduct an additional sequence-length balancing operation on the samples. This ensures that the average lengths of the positive and negative samples are close to each other.

### v2

@@ Abstract
In recent years, the field of vision-language model pre-training has experienced rapid advancements, driven primarily by the continuous enhancement of textual capabilities in large language models. However, existing training paradigms for multimodal large language models heavily rely on high-quality image-text pairs. As models and data scales grow exponentially, the availability of such meticulously curated data has become increasingly scarce and saturated, thereby severely limiting further advancements in this domain. This study investigates scalable caption generation techniques for vision-language model pre-training and demonstrates that large-scale low-hallucination synthetic captions can serve dual purposes: 1) acting as a viable alternative to real-world data for pre-training paradigms and 2) achieving superior performance enhancement when integrated into vision-language models through empirical validation. This paper presents three key contributions: 1) a novel pipeline for generating high-quality, low-hallucination, and knowledge-rich synthetic captions. Our continuous DPO methodology yields remarkable results in reducing hallucinations. Specifically, the non-hallucination caption rate on a held-out test set increases from $48.3\%$ to $77.9\%$ for a 7B-size model. 2) Comprehensive empirical validation reveals that our synthetic captions confer superior pre-training advantages over their counterparts. Across 15 vision language tasks, the model trained with our data achieves a significant performance gain of at least $6.2\%$ compared to identical images with alt-text. In 20 common cognitive domains, the model trained with our data outperforms the alt-text data by at least $7.5\%$. Meanwhile, it also offers considerable support in the text-to-image domain. With the support of our dataset, the FID score is reduced by $17.1$ on a real-world validation benchmark and $13.3$ on the MSCOCO validation benchmark. 3) We will release Hunyuan-Recap100M, a low-hallucination and knowledge-intensive synthetic caption dataset, which will enable breakthroughs in vision language pre-training, cross-modal generation, and related applications.
@@ 1 Introduction
With the enormous power of large language models, remarkable performance gains have recently been achieved in a variety of tasks in natural language processing (NLP), computer vision (CV), and also in cross-modal fields (Brown et al., 2020; Chung et al., 2022; Chowdhery et al., 2022; Touvron et al., 2023; Dosovitskiy et al., 2021; Alayrac et al., 2022; Liu et al., 2023; Achiam et al., 2023; Wang et al., 2024; Hurst et al., 2024). The Large Language Model (LLM) and the Vision Language Model (VLM) are usually equipped with Transformer (Vaswani et al., 2017) as the backbone and then pre-trained with a tremendous amount of unlabeled data. The strong representation ability of the model, the massive amount of data, and the effective means of training make the foundation models powerful for successfully solving the tasks of vision and language.
@@
While text data has experienced exponential growth and fuels large language models, vision language model pre-training confronts a unique data bottleneck: high-quality text-image alignments remain scarce despite their critical role in training vision language models. Previous approaches to building large-scale multimodal datasets rely on Web-based alt-text extraction (Sharma et al., 2018; Changpinyo et al., 2021; Ordonez et al., 2011; Schuhmann et al., 2021, 2022; Byeon et al., 2022; Gadre et al., 2023; Liu et al., 2022; Gu et al., 2022). The alt-text data undergo a lengthy processing pipeline, involving stages such as text cleaning, image filtering, and cross-modal relevance validation, ultimately producing image-text pairs. Although significant data loss occurs at each processing stage and even with these efforts, the final output still suffers from inherent limitations such as low informativeness and partial misalignment between visual content and the corresponding descriptions. As naturally occurring high-quality image-text pairs are increasingly exhausted, traditional extraction methods experience considerable declines in output quality when scaled up, leading to significantly diminishing marginal returns in the training of VLMs. Therefore, novel techniques are urgently needed to produce large-scale, high-quality multimodal datasets.
@@
Recently, several exploratory investigations have been carried out in multi-modal synthetic data, which have, to some extent, demonstrated their positive impact on various cross-modal tasks (Li et al., 2023a; Betker et al., 2023; Yu et al., 2024; Li et al., 2024b). However, there are significant defects in all of them, mainly reflected in the presence of hallucinations and low information density. As shown in Figure 1, there are a large number of hallucinations (labeled in red) in the previous synthetic captions, and the captions are rather monotonous, lacking informative expressions. Therefore, in this paper, we focus on addressing two issues: 1. How to establish an effective pipeline for generating low-hallucination synthetic captions? 2. How can more valuable information be injected into synthetic captions to increase the information density of synthetic data?
@@
How to establish an effective pipeline for generating low-hallucination synthetic captions? VLMs enable efficient large-scale caption synthesis, but persistent hallucination effects constrain their reliability (Bai et al., 2024). As shown in Figure 1, Recap-DataComp-1B (Li et al., 2024b) contains a significant amount of hallucinatory information. Excess hallucinatory information in the training data is ultimately encoded in the final model, hindering the learning of correct patterns. The severity of this problem is quantitatively demonstrated in Section 4. Given the propensity for models trained on imperfect data to produce unfounded descriptions, effectively combating the generation of hallucinatory captions remains a paramount and often unavoidable challenge in caption synthesis. To address the crucial issue of hallucination in synthesized captions, we propose employing the Direct Preference Optimization (DPO) method. This approach aims to mitigate hallucinations by increasing the model’s confidence in generating descriptions of real visual content while simultaneously suppressing the confidence associated with hallucinatory visual elements. Furthermore, within the context of the synthetic caption task, we observe that DPO adheres to a significant scaling law within a specific data-size range. Notably, upon saturation of this scaling behavior, resampling of the data can further enhance the quality of the synthetic captions. As demonstrated in Table 1, applying our proposed method (more detailed in Section 3) to the same dataset yields a significant improvement in the non-hallucination rate (evaluated using metrics described in Section 4.1). Specifically, our approach increases the non-hallucination rate from 29.7% to 73.3%, representing a substantial enhancement compared to the baseline approach used in Recap-DataComp-1B (Li et al., 2024b).
@@
Table 1: A comparison between existing synthetic caption datasets and our proposed dataset in terms of hallucinations. Here, "Length" represents the average length of the caption, "Details" represents the average number of visual details that appear in each caption, "None Hallucinations" indicates the proportion of non-hallucinatory data in the overall data, and "hallucination rate" represents the proportion of hallucinations among all details. A total of 1,000 data samples were taken for validation in each dataset.
How can more valuable information be injected into synthetic captions to increase the information density of synthetic data? VLM outputs often prioritize visual saliency, resulting in captions that are generally correct but informatively sparse, lacking deeper context and specific details derived from external or implicit knowledge. To overcome this limitation and produce synthetic captions with high information density, our approach involves training a VLM to effectively synthesize both external textual knowledge (e.g. alt-text, metadata) and its inherent world knowledge. Specifically, we used a powerful VLM, GPT-4o, guided by a meticulously designed prompt, to create the initial set of labeled data. A key aspect of this data generation process is the focused extraction and integration of vital information embedded in external text sources, which serves as a readily available source of descriptive and often factual information about images. For example, particular attention is paid to identifying and incorporating specific entity names or critical attributes within the alt-text, illustrated here by the case of oxalic acid in Figure 1. Subsequently, the generated data undergoes a rigorous manual post-processing step to rectify obvious errors, confirm the inclusion of essential key information targeted during the generation phase, and effectively filter out noise or irrelevant details introduced during the initial knowledge injection. Finally, an open-source VLM, Qwen2-VL-7B, is fine-tuned on this carefully curated and information-rich dataset.
Building upon the meticulously designed data generation and model training pipeline described previously, we developed and trained a specialized recaption model. We then proceeded to generate a large-scale synthetic dataset, Hunyuan-Recap100M, comprising 100 million image-text pairs. Compared to models pre-trained using raw alt-text data or conventional synthetic captions, VLMs trained with the Hunyuan-Recap100M dataset exhibit substantially improved overall performance. This significant performance gain, particularly pronounced in tasks requiring fine-grained visual perception, highlights the critical impact of training on synthetic data that not only offers high information density (as discussed previously), but also incorporates significantly reduced hallucinatory content, thereby enhancing the robustness and factual grounding of the downstream models. Empirical tests show models trained with our data gain at least 6.2% on 15 VL tasks and 6.8% on 20 cognitive domains compared to alternatives. This data also benefits text-to-image generation, reducing FID on validation benchmarks by up to 17.1.
Our contributions are summarized as follows:
1. High-Quality Synthetic Data Generation: We develop a novel pipeline leveraging specialized methods to generate synthetic captions that are both significantly low in hallucination (improving the hallucination-free rate from $48.3\%$ to a notable $77.9\%$) and rich in knowledge.
2. Large-Scale Dataset and Demonstrated Impact: We construct and will release Hunyuan-Recap100M, a substantial 100 million image-text dataset generated by our pipeline. We demonstrate that pre-training vision-language models with Hunyuan-Recap100M leads to substantially improved overall performance compared to existing synthetic captions, particularly enhancing perceptual capabilities, thus providing a valuable resource to the community.
@@ 2 Related Work
Vision-Language Model. Vision Language Models (VLMs) can be broadly classified into two categories. The first utilizes dual encoders, typically comprising a vision and a text encoder. These models are trained contrastively to learn aligned cross-modal representations and are highly effective in tasks such as image-text retrieval and zero-shot classification. Prominent examples include CLIP (Radford et al., 2021), FLIP (Li et al., 2023b), and SigLIP (Zhai et al., 2023b). The second category focuses on integrating visual information with a Large Language Model (LLM). Driven by the remarkable success of LLMs (Brown et al., 2020; OpenAI, 2023; Grattafiori et al., 2024; Guo et al., 2025), endowing these models with visual understanding has become a significant research direction. Flamingo (Alayrac et al., 2022) pioneered large-scale VLM research by aligning a visual encoder with an LLM. LLaVA (Liu et al., 2023) further established a methodology for training large multimodal models through instruction tuning. Subsequent efforts have continued to refine specific modules and optimize the overall performance of VLMs (Wang et al., 2024; Wu et al., 2024; McKinzie et al., 2024; Chen et al., 2024b; Liu et al., 2024). Although current LLM-based vision language models demonstrate significantly greater capabilities compared to earlier multimodal models (Zeng et al., 2021; Singh et al., 2021; Diao et al., 2022; Wang et al., 2022; Zhang et al., 2023a), they fundamentally share a core dependency with dual encoder VLMs (the first category discussed previously): the need for large-scale image-text pairs for training. This reliance presents a growing challenge, however, as the supply of high-quality image-text pair data is finite (Villalobos et al., 2022) and its expansion struggles to keep pace with the increasing scale of models. To overcome this data bottleneck, synthetic data offers a promising potential solution.
Synthetic data. Building on its success in enhancing large language models (Zhang et al., 2024; Shao et al., 2024; Zhu et al., 2024; Yang et al., 2024), synthetic data generation is increasingly being explored within the multimodal domain. Some methods, such as BLIP-2 (Li et al., 2023a), utilize image captioning models to synthesize short captions as substitutes for alt-text, with CapFusion (Yu et al., 2024) leveraging LLMs for refinement. However, captions generated by these approaches often remain simplistic, providing minimal added information value beyond the original alt-text. Recap-DataComp-1B (Li et al., 2024b) attempts to transform alt-text using existing vision language models, but faces the significant challenge of hallucination inherent in current VLMs. Other works focusing on dense captioning, including Allava (Chen et al., 2024a) and Dense-Fusion-1M (Li et al., 2024c), generate localized descriptions, but the resulting datasets can also suffer from notable hallucinations. These limitations in existing synthetic VLM data highlight the need for more sophisticated generation techniques. In summary, generating high-quality multimodal synthetic data faces challenges related to knowledge enrichment and hallucination control. Although VLM hallucination is a studied topic (Bai et al., 2024), a comprehensive synthetic visual captioning method capable of simultaneously producing diverse and rich content with minimal hallucination is still needed.
Our approach makes substantial progress in the generation of synthetic visual captions with improved knowledge injection and significantly reduced hallucination. This yields a caption dataset whose efficacy is demonstrated for large-scale VLM training, offering a viable alternative that overcomes the constraints of limited real data and the quality challenges of scaling traditional datasets.
@@ 3 Methodology
In this section, we present our synthetic data pipeline, including continuous DPO aimed at reducing hallucinations, knowledge-enriching SFT to enhance knowledge content, and details of the model training pipeline and hyper-parameters.
@@ 3.1 Continuous DPO
Reinforcement learning has been reported to be an effective method of mitigating hallucinations in large language models (Zhang et al., 2023b; Huang et al., 2025). Direct Preference Optimization (DPO) (Rafailov et al., 2023) emerges as a reinforcement learning based fine-tuning paradigm for LLM that directly optimizes preference data, effectively mitigating harmful content generation, such as hallucinations. Consequently, the application of DPO to mitigate hallucinations in image captions appears to be a natural and logical approach. The optimization objectives of the DPO algorithm are shown in Equation 1 and Equation 2. As noted in Rafailov et al. (2023), the loss function gradient $\mathcal{L}_{\text{DPO}}$ increases the probability of preferred completions $y_{w}$ while simultaneously reducing the probability of dispreferred completions $y_{l}$. Intuitively, by maximizing the probability of samples free of hallucinations and minimizing the probability of samples containing hallucinations, our recaption model can generate captions that are as free from uncertainty and potential hallucinations as possible.
@@
However, in our task, standard DPO does not exhibit continuous performance improvement with respect to data scale. Specifically, after initial rapid performance gains, its performance plateaus, as illustrated in Figure 4. Examining the DPO objective (Equation 2), we see that the training examples are weighted based on the implicit reward derived from the reference model, which reflects the preference strength between chosen and rejected completions. As the scale of the DPO dataset increases, the implicit reward signal, derived from the fixed reference model, does not change significantly. Consequently, the optimization plateaus as the model converges to a local optimum defined by this relatively static signal. To address this limitation, we propose Continuous DPO (CDPO), a modification of the standard DPO algorithm. In CDPO, when performance plateaus, we iteratively update the reference model (e.g., using the current policy), resample preference data using the updated policy, which in turn yields a new implicit reward signal, and resume optimization. Furthermore, during CDPO training, to prevent the sample length from acting as an unintended learning signal (e.g. if longer completions are more likely to contain hallucinations), we apply a sequence-length balancing operation to the preference pairs. This ensures that the average lengths of the preferred and dispreferred completions within the dataset are closely matched.